Range Balancing: 13 February 2011 UTC

Observer(s): John Hill, Andrew Rakich(Tucson), Doug Miller (home)

Telescope Operator: Geno Bechetti (LBTO)

SW Support: (Tucson)

Telescope Support: (LBTO)

Instrument Support: (LBTO)

Summary

Photometric, with seeing between 0.3 and 0.5 arcsec all night long. Temperature is a balmy 6 degC. Can we be flogged for taking pointing data under these conditions?

TCS people had the first half of the night for testing BP9 binocular stuff. We successfully sent a synchronized acquire preset that produced stars on both sides, although it was lots of work to get there.

During the second half, we tested Source Extractor active optics. Collected 52 stars for pointing data. Measured field aberrations at low elevations.

Details

TCS Testing

01:00 Paul and Tom finishing up tests of swing arms and dynamic balancing.

02:17 Adjusting pointing on WT10_229, M5_0202, and M5_0050. SX IE=-72 CA=0, DX IE=4 CA=-12

+20 Tilt combines with -20 DETXY offset.

Applied -2000 nm Z7 to DX M2.

At EL=46, DX IE=+37 CA=-45

04:00 PCS is doing crazy stuff with the tip-tilts when we send an offset.

04:11 Consecutive absolute synchronized offsets with 0,0 DETXY are causing the telescope to make ~1 arcminute excursions in tracking. very peculiar!

04:30 We patch PCS to eliminate the feedback of actual PCS positions. Now the tip-tilts are equal and opposite between the two sides.

04:39 switch from binocular to monocular (need to zero the offsets on both sides)

04:50 Monocular Pointing Adjustment at EL=68
SX IE=-75 CA=+10
DX IE=+3 CA=0

05:00 Can see what seems to be vibrations in 2 sec star trails being collected by GCS in acquire mode while telescope is not tracking. (this data wasn't saved.) Further discussion with D. Ashby suggests that what we saw was not vibration, but hunting of the servo in low-gain mode while it was holding.

Rebooted IRTC windows PC.

05:24 Install a PCS patch that isn't hammering LSS with too many events.

05:30 Signs are clearly reversed for the optics tip-tilts. John demonstrated this by find the stars after applying -2x the tip-tilt corrections manually. Reversed the signs in lbt.conf.

05:50 Changed the guide rate in AIP_L.cfg for BP9 from 0.4 to 0.04 Hz.

06:00 Synchronous Active Preset to BS9126. Acquisition offset wasn't very good. Guide corrections are OK. Then telescope stopped tracking because of polys late by several seconds. PCS is causing this somehow?

06:30 Patch PCS to eliminate the Pause/Resume guiding calls to GCS when PSF tip-tilts are updated. This eliminates the very late polys, but we don't understand the actual cause. There is still a problem that GCS isn't waiting for the PSF to implement/finish the acquisition offset before taking the first guide image.

Things that are unresolved:
- PCS late polys
- GCS communication confuses PCS
- tip-tilts seem not to match LBC results

Sextractor Active Optics Tests

7:30 Going to try Doug's Sextractor active optics.

7:35 changed to BP7. Corrected pointing on M5_0347. Active preset to BS9138. IE -75 CA +10

8:10 GS 0 BS91238. Doug running IDL from home. Sxtr first diverged, then after a fix converged nicely. Seeing spotting down to 0.26".

8:12 Active preset to Stone_G, GS1.

LFBG Pointing Data

8:28 turning on pointing logging. 20110213_sx.log, 20110213_dx_log.

Using /home/LBTO/iraf/gopoint_agw.cl (This version is hardwired for left side.) Tonight we are using it to send GUIDE presets (hardwired). After preset completes, we sleep for 10 seconds, stop guiding, and then take a 1 second exposure with the guide camera.

Stopped and power cycled AGW3 trying to solve the AzCam problem.

Pointing images are in /OldRepository/AGW_Data/20110213. These images are guided, so the images are only needed to provide time entries into the pointing log. Still using adjusted parameters in AIP_L.cfg for bright pointing stars.

10:45 Completed 52 stars.

Field Aberration Data at low elevation

10:55 preset to BS9177 and run field_collect.

11:10 Note that wfsc 91 represents a case where the IDL routine collimated 3 times, got a final wavfront of 373 nm, and then went off to a guide star without announcing it was collimated or converging to the normal limit.

11:11 note that the interrupt in the field_collect routine came about because "ctrl C" in Windows means "copy", but it means something different for IDL.

11:25 GS20 (#6 in sequence) saturated the WF camera, restarted at #7. We removed it from the list (BS9177_field.list) later.

12:12 Completed first pass through the list at EL=41 deg. Started second pass.

13:17. Got to star 12 or 13 on second pass, closing for dawn.

-- AndrewRakich - 13 Feb 2011
Topic revision: r11 - 25 Apr 2011, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback